OpenStudy (anonymous):

Please help me?? My answer is A If f(x) = 3x + 1 and g(x) = x2 + 2x + 1, what is f(g(x)) ? (1 points) Answer 3x2 + 2x + 2 3x2 + 6x + 4 x2 + 5x + 2 x2 - 5x + 2

OpenStudy (anonymous):

f(g(x)) = 3(x2 + 2x + 1) + 1 , = 3x2 + 6x + 3 + 1 , = 3x2 + 6x +4